# Deductive Beam Search: Decoding Deducible Rationale for Chain-of-Thought Reasoning

### v2

@@ 6 Analysis
Table 4: Verification ability comparison of our verifier and ChatGPT. We ask both models to rank among the gold reasoning steps and the synthesized reasoning steps. The evaluation focuses on determining the position these models assigned to the gold steps.
@@
6.1 Verifier Analysis
We comprehensively analyze the verifiability of our proposed verifier. We analyze performance (Table 4), score distribution (Figure 3), and robustness on different model scales (Figure 4).
Empirical Study. For empirical experiments, we test how our verifier will score the gold reasoning steps compared to synthesized reasoning steps. We randomly sample 500 context-step pairs from the test set of the GSM8K dataset as gold reasoning steps. For each pair, we synthesize nine reasoning steps using Llama2-7b as inferior reasoning steps. The verifier’s task is to rank them, and the performance is evaluated by four metrics, namely, Mean Reciprocal Rank (MRR), HITS@1/3/5. MRR evaluates the average rank of the gold reasoning step, and the HITS metrics reflect whether the gold reasoning step will be chosen under beam size setting 1/3/5. To compare with the self-evaluation pattern, we ask ChatGPT to rank these reasoning steps rather than predict scores, leveraging its inherent reranking capabilities (Ma et al., 2023). The results are listed in Table 4. Our verifier outperforms ChatGPT across all metrics, evidencing its capability. Notably, our verifier correctly identifies 86% of the gold reasoning steps within the top 5 positions out of 10 samples, affirming the deducibility of the reasoning paths decoded under $m=5,n=10$.
@@ 6.1 Verifier Analysis
Figure 3: Distributions of language model and verifier scores on correct/incorrect reasoning paths. The dashed lines represent the median. Our verifier differentiates the scores attributed to correct and incorrect reasoning paths.
Figure 4: Accuracy under different deductive score thresholds on greedy decoding results from Llama2-7b and ChatGPT. The dashed line shows the polynomial fitting the scattered data. The near-parallel of these lines suggests the robustness of our verifier’s performance.
@@
Robustness. To ensure that our verifier can equally verify reasoning steps generated by different models, we visualize the accuracy under different deductive score thresholds for Llama2-7b and ChatGPT, as depicted in Figure 4. The lines in the figure represent polynomial fits of the data. Their near-parallel alignment suggests the robustness of performance improvement across these models as the threshold increases. Intriguingly, these lines also offer insights into the inductive reasoning capabilities of the two models. Ideally, the accuracy approaches zero at the deductive score zero. However, the observed non-zero accuracy suggests the models’ inductive reasoning capabilities. Although the reasoning process might not align with the deductive reasoning paradigm, the models can still arrive at correct conclusions, likely by intuitively skipping over specific reasoning steps, which is the act of inductive reasoning.
@@ 6.2 Cost Analysis
The cost of sampling multiple times is enormous. We analyze the cost of our methods under different settings and compare them with the baselines. Specifically, we compare our approach with greedy decoding, self-consistency, and self-evaluated decoding strategies (Xie et al., 2023). The results are presented in Table 3 and Table 5. Our analysis reveals that greedy decoding is the most token-economic, as it does not involve any form of sampling, but its performance lags. When $m$ is constrained to 1, the token generation is minimized even further than that required by self-consistency strategies. Still, our method demonstrates higher accuracy. Moreover, DBS proves more effective and token-efficient under the same beam size than those leveraging LLMs’ self-evaluation capabilities.
Table 5: Cost Analysis of our method and other baseline methods. We report output tokens for each method. We use Llama2-7b as the backbone for all methods. $n$ represents sampling times of the method, and $m$ represents beam size.
Method Accuracy Tokens
Greedy 21.99 267,462
Self-consistency, $n=10$ 28.05 1,984,921
DBS , $m=1$, $n=10$ 29.34 1,414,435
DBS , $m=5$, $n=10$ 31.16 4,042,053
Table 6: Different prompt settings on commonsense reasoning tasks. We use Llama2-7b as the backbone for all methods. For methods with multiple reasoning chains, we take 10 reasoning paths for self-consistency.
@@
   - w. recall prompt 65.80 69.45
@@
Question Greedy Ours
James decides to run 3 sprints 3 times a week. He runs 60 meters each sprint. How many total meters does he run a week?
He runs 3 times a week, which is equal to 3 x 7 = 21 sprints. 0.0102 The distance he runs each sprint is 60 meters, so the total distance he runs is 60 x 21 = 1320 meters. 0.4963 Final Answer:1320
Each sprint is 60 meters. 0.8270 So he runs 60 x 3 = 180 meters each day. 0.9643 He runs each sprint 3 times a week. 0.9983 Therefore, he runs a total of 180 x 3 = 540 meters per week. 0.9463 Final Answer:540
Toulouse has twice as many sheep as Charleston. Charleston has 4 times as many sheep as Seattle. How many sheep do Toulouse, Charleston, and Seattle have together if Seattle has 20 sheep?
Toulouse has twice as many sheep as Charleston, so Toulouse has 2 x 4 = 8 sheep. 0.3149 Charleston has 4 times as many sheep as Seattle, so Charleston has 4 x 20 = 80 sheep. 0.9877 So Toulouse has 8 + 80 = 88 sheep. 0.3923 Seattle has 20 sheep. 0.7479 Final Answer:88 + 20 + 88 = 196
Let’s first find out how many sheep Charleston has. 0.9447 Charleston has 4 times as many sheep as Seattle. Seattle has 20 sheep. Charleston has 20 x 4 = 80 sheep. 0.9912 Now, let’s find out how many sheep Toulouse has. Toulouse has twice as many sheep as Charleston. Charleston has 80 sheep. Toulouse has 80 x 2 = 160 sheep. 0.9736 Finally, let’s find the total number of sheep Toulouse, Charleston, and Seattle have together. Add the number of sheep Toulouse, Charleston, and Seattle have. 0.9981 Toulouse has 160 sheep, Charleston has 80 sheep, and Seattle has 20 sheep. So the total number of sheep they have together is 160 + 80 + 20 = 260 sheep. 0.9924 Final Answer:260
6.3 Commonsense Reasoning Task
Unlike arithmetic or symbolic reasoning tasks, commonsense reasoning tasks leverage the model’s intrinsic commonsense. Thus, we introduce a recall prompt, which asks the model to recall the necessary commonsense via greedy decoding, followed by DBS navigating the reasoning over the elicited contexts. We compare this setting to standard settings and baselines, and the results are shown in Table 6. The performance varies from different tasks, attributed to the distinct nature of the tasks. Concretely, StrategyQA typically requires a 2-3 step knowledge recall followed by a single reasoning step, whereas CSQA demands a more extensive 3-5 step recall process alongside multiple reasoning steps. Our findings suggest that the recall prompt is more suitable for tasks demanding comprehensive commonsense and multi-step reasoning.
@@ 7 Related Work
Self-Evaluation. Recent research on reducing reasoning errors is inclined to follow the self-verify-then-correct pattern (Dhuliawala et al., 2023; Weng et al., 2023; Zhang et al., 2023b; Ling et al., 2023; Miao et al., 2023). They design different procedures and prompts to achieve better performance. Taking two typical approaches as examples, Dhuliawala et al. (2023) design a chain-of-verification procedure to verify facts from its outputs, and Miao et al. (2023) ask LLM to detect errors in their step-by-step rationales. However, recent works (Huang et al., 2023; Hong et al., 2023) have pinpointed a critical limitation of LLMs in self-correction during reasoning tasks. Their findings suggest that LLMs indiscriminately alter reasoning steps without external feedback, irrespective of their initial accuracy.
Decoding Strategies. Conventional decoding strategies include greedy decoding (Teller, 2000), which selects tokens with the highest probabilities, and beam search (Graves, 2012), which stores candidate beams for future prediction. In the era of LLMs, these decoding strategies are implemented at a more coarse-grained level, especially on reasoning tasks involving multiple steps in the solving process. They decompose the reasoning process into steps (Khot et al., 2022) and apply decoding or search algorithms (Yao et al., 2023; Xie et al., 2023).
@@ 8 Conclusions
In this paper, we aim to eliminate errors in intermediate reasoning steps in CoT reasoning. To this end, we propose Deductive Beam Search that integrates CoT with step-wise beam search and scores each reasoning step with a deductive verifier, which verifies whether the reasoning step is a logical consequence. Beam search explores by sampling potential reasoning steps, while the verifier exploits by selecting the most deducible steps. To train such a verifier, we propose a scalable and labor-free data construction method. It initiates by heuristically introducing errors into gold reasoning steps and enhances the diversity and difficulty of training data by synthesizing hard negatives through the verifier trained on those typical wrong steps. Extensive experiments show our method’s effectiveness across various model scales and diverse reasoning tasks without changing the standard CoT paradigm and parameters of LLMs. Further analysis proves the verifiability and robustness endowed by our verifier, thereby significantly improving the deducibility of the generated reasoning paths.
@@ Appendix A Extended Experiments
Comparison with more baselines.
We compare more baselines on GSM8K, using Llama2-7b as the backbone. All methods are implemented with the same prompt for reasoning step generation. Our method outperforms this baseline on both single chain setting and multiple chain setting. Notably, these methods utilize self-evaluation, indicating that the self-evaluation capability is weak on a relatively small language model.
@@ B.2 Training Details
Table 8: Comparison of baselines focusing on correcting intermediate errors. DV represent Ling et al. (2023), and SelfEval represents Xie et al. (2023).
Method Accu #Rationales
Greedy 21.99 1
DV 10.46 1
SelfEval 22.76 1
Ours 31.16 1
[…]
SelfEval + SC 24.18 10
Ours + SC 32.07 10
@@ B.3 Prompts
Table 10: Prompt for StrategyQA.
@@
Table 11: Prompt for CSQA.
@@
Table 12: Prompt for Coin.
@@
Table 13: Prompt for StrategyQA with prompt.
@@
Table 14: Prompt for CSQA with recalling commonsense first.

### v3

@@ 6 Analysis
6.1 Verifier Analysis
@@
Table 4: Verification ability comparison of our verifier and ChatGPT.
We comprehensively analyze the verifiability of our proposed verifier on performance (Table 4) and score distribution (Figure 3).
Empirical Study.
For empirical experiments, we test how our verifier will score the gold reasoning steps compared to synthesized reasoning steps. We randomly sample 500 context-step pairs from the test set of the GSM8K dataset as gold reasoning steps. For each pair, we synthesize nine reasoning steps using Llama2-7b as inferior reasoning steps. The verifier’s task is to rank them, and the performance is evaluated by four metrics, namely, Mean Reciprocal Rank (MRR), HITS@1/3/5. MRR evaluates the average rank of the gold reasoning step, and the HITS metrics reflect whether the gold reasoning step will be chosen under beam size setting 1/3/5. To compare with the self-evaluation pattern, we ask ChatGPT to rank these reasoning steps rather than predict scores, leveraging its inherent reranking capabilities (Ma et al., 2023). The results are listed in Table 4. Our verifier outperforms ChatGPT across all metrics, evidencing its capability. Notably, our verifier correctly identifies 86% of the gold reasoning steps within the top 5 positions out of 10 samples, affirming the deducibility of the reasoning paths decoded under $m=5,n=10$.
@@
Figure 3: Distributions of language model and verifier scores on reasoning paths.
Method Accuracy Tokens Greedy 21.99 267,462 Self-consistency, $n=10$ 28.05 1,984,921 DBS , $m=1$, $n=10$ 29.34 1,414,435 DBS , $m=5$, $n=10$ 31.16 4,042,053 Table 5: Cost Analysis of our method and other baseline methods. $m$ represents beam size and $n$ represents sampling times.
@@ 6.2 Cost Analysis
The cost of sampling multiple times is enormous. We analyze the cost of our methods under different settings and compare them with the baselines. Specifically, we compare our approach with greedy decoding, self-consistency, SelfEval (Xie et al., 2023), and DV (Ling et al., 2024). The results are presented in Table 3 and Table 5. Our analysis reveals that greedy decoding is the most token-economic, as it does not involve any form of sampling, but its performance lags. When $m$ is constrained to 1, the token generation is minimized even further than that required by self-consistency strategies. Still, our method demonstrates higher accuracy. Moreover, DBS proves more effective and token-efficient under the same beam size than those leveraging LLMs’ self-evaluation capabilities.
6.3 Commonsense Reasoning Task
@@
- w. recall prompt 65.80 69.45
Table 6: Different prompt settings on commonsense reasoning tasks. We use Llama2-7b as the backbone for all methods.
Unlike arithmetic or symbolic reasoning tasks, commonsense reasoning tasks leverage the model’s intrinsic commonsense. Thus, we introduce a prompt strategy, which asks the model to recall the necessary commonsense via greedy decoding, followed by DBS navigating the reasoning over the elicited contexts. We compare this setting to standard settings and baselines, and the results are shown in Table 6.
Question Greedy Ours
James decides to run 3 sprints 3 times a week. He runs 60 meters each sprint. How many total meters does he run a week? He runs 3 times a week, which is equal to 3 x 7 = 21 sprints. 0.0102 The distance he runs each sprint is 60 meters, so the total distance is 60 x 21 = 1320 meters. 0.4963 Final Answer:1320 Each sprint is 60 meters. 0.8270 So he runs 60 x 3 = 180 meters each day. 0.9643 He runs each sprint 3 times a week. 0.9983 Therefore, he runs a total of 180 x 3 = 540 meters per week. 0.9463 Final Answer:540
Toulouse has twice as many sheep as Charleston. Charleston has 4 times as many sheep as Seattle. How many sheep do Toulouse, Charleston, and Seattle have together if Seattle has 20 sheep? Toulouse has twice as many sheep as Charleston, so Toulouse has 2 x 4 = 8 sheep. 0.3149 Charleston has 4 times as many sheep as Seattle, so Charleston has 4 x 20 = 80 sheep. 0.9877 So Toulouse has 8 + 80 = 88 sheep. 0.3923 Seattle has 20 sheep. 0.7479 Final Answer:88 + 20 + 88 = 196 Charleston has 4 times as many sheep as Seattle. Charleston has 20 x 4 = 80 sheep. 0.9912 Toulouse has twice as many sheep as Charleston. Charleston has 80 sheep. Toulouse has 80 x 2 = 160 sheep. 0.9736 Finally, add the number of sheep Toulouse, Charleston, and Seattle have. 0.9981 So the total number of sheep they have together is 160 + 80 + 20 = 260 sheep. 0.9924 Final Answer:260
@@
The performance varies from different tasks, attributed to the distinct nature of the tasks. Concretely, StrategyQA typically requires a 2-3 step knowledge recall followed by a single reasoning step, whereas CSQA demands a 3-5 step recall process alongside multiple reasoning steps. Our findings suggest that the recall prompt is more suitable for tasks demanding multi-step reasoning.
@@ 7 Related Work
Self-Evaluation. Recent research on reducing reasoning errors is inclined to follow the self-verify-then-correct pattern (Dhuliawala et al., 2023; Weng et al., 2022; Zhang et al., 2023b; Ling et al., 2024; Miao et al., 2023). They design different procedures and prompts to achieve better performance. Taking two typical approaches as examples, Dhuliawala et al. (2023) design a chain-of-verification procedure to verify facts from its outputs, and Miao et al. (2023) ask LLM to detect errors in their step-by-step rationales. However, recent works (Huang et al., 2023; Hong et al., 2023; Xie et al., 2024) have pinpointed a critical limitation of LLMs in self-correction during reasoning tasks. Their findings suggest that LLMs indiscriminately alter reasoning steps without external feedback, irrespective of their initial accuracy.
Decoding Strategies. Conventional decoding strategies include greedy decoding (Teller, 2000), which selects tokens with the highest probabilities, and beam search (Graves, 2012), which stores candidate beams for future prediction. In the era of LLMs, these decoding strategies are implemented at a more coarse-grained level, especially on reasoning tasks involving multiple steps. They decompose the reasoning process into steps (Khot et al., 2022) and apply decoding or search algorithms (Yao et al., 2023; Xie et al., 2023).
@@ 8 Conclusions
In this paper, we aim to eliminate errors in intermediate reasoning steps in CoT reasoning, making it more reliable. To this end, we propose Deductive Beam Search that integrates CoT with step-wise beam search and scores each reasoning step with a deductive verifier, which verifies whether the reasoning step is a logical consequence. Beam search explores by sampling potential reasoning steps, while the verifier exploits by selecting the most deducible steps. To train such a verifier, we propose a scalable and labor-free data construction method. It initiates by heuristically introducing errors into gold reasoning steps and enhances the diversity and difficulty of training data by synthesizing hard negatives through the verifier trained on those typical wrong steps. Extensive experiments show our method’s effectiveness across various model scales and diverse reasoning tasks without changing the standard CoT paradigm and parameters of LLMs. Further analysis proves the verifiability and robustness endowed by our verifier, thereby significantly improving the deducibility of the generated reasoning paths.
Acknowledgments
The authors would like to thank Jiahao Shi and Yikai Zhang from Fudan University, and Yifei Li from the Ohio State University, as well as the anonymous reviewers for their valuable comments.
@@ Appendix A Extended Experiments
Figure 4: Accuracy under different beam sizes on different models.
Figure 5: Accuracy under different deductive score thresholds on greedy decoding results.
@@
Verifier Robustness.
To ensure that our verifier can equally verify reasoning steps generated by different models, we visualize the accuracy under different deductive score thresholds for Llama2-7b and ChatGPT, as depicted in Figure 5. The lines in the figure represent polynomial fits of the data. Their near-parallel alignment suggests the robustness of performance improvement across these models as the threshold increases. Intriguingly, these lines also offer insights into the inductive reasoning capabilities of the two models.
Figure 6: Accuracy under different reasoning step length.
Ideally, the accuracy approaches zero at the deductive score zero. However, the observed non-zero accuracy suggests the models’ inductive reasoning capabilities. Although the reasoning process might not align with the deductive reasoning paradigm, LMs can still arrive at correct conclusions, likely by intuitively skipping over specific reasoning steps, which is the act of inductive reasoning.
DBS Robustness. To demonstrate that DBS can sustain its accuracy as the length of the reasoning steps increases, we conducted experiments to analyze the impact of reasoning step length on the final outcome’s accuracy. The results, presented in Figure 6, indicate that DBS consistently achieves higher accuracy, even when the reasoning process extends to 15 steps.
@@ B.3 Prompts
Table 8: Prompt for arithmetic reasoning tasks.
@@
Table 9: Prompt for StrategyQA.
@@
Table 10: Prompt for CSQA.
@@
Table 11: Prompt for Coin.
@@
Table 12: Prompt for StrategyQA with prompt.
@@
Table 13: Prompt for CSQA with recalling commonsense first.